A social psychologist wants to test if extroversion is significantly related to a person’s general health. He gets extroversion scores and a measure of general health from the participants’ medical providers. The data are listed below. Higher health extroversion scores mean greater extroversion, while high health scores = good health.  

Complete the appropriate statistical test and hypothesis testing using the 4 step approach. State your conclusion in APA format. Alpha level = .05.

Participant

extrov X

health Y

X squared Y squared (X) (Y)
A

101

33 3333
B 93 40 3720
C 85 24 2040
D 111 46 5106
E 104 30 3120
F 89 28 2492
Total

Using the data from the prior correlation question (extroversion and health),

Find the regression line for the data.

a = ____       b = ______        equation:   _____________________
Answer #1

solution given data; h = 6 Ex=583 Ey=201 Exy - 19811 Ssx = 282 hiss = 310 2.05 ssay {(xi-xj Cy; -ý exy-nX 7 T = Ex 583 -97.16ų = Ey 201 6 = 33.5 Y E = 33.5 ssly = 19811 – 33.5X97.1667 X6 =19811 - 19530.5067 SSXY = 280.4933 to test we want Hoë there i3 test stastistics a » &n-2 -tn-2 Vlore 0.9486/6-2 = 1.8972 599912 1 -0.94862 0:3/64 P-value =0.002 the sitical tn-2, 4/2 = t